What is t ethanoic acid in question 6 to find K, and pK, for the ethanoate ion (CH,COO). Hint: The ethanoate ion is the conjugate base of ethanoic acid he value of Kw at 25°C? Use this value and the acid dissociation constant given for
Answer #1

The value of Kw at 250c is 1*10-14 .

The reaction is

CH3COOH +H2O = CH3COO- + H3O+  

We know, Kw= [H+] [OH-] =1*10-14

Kw= Ka*Kb

or, Kb= Kw/Ka

or, -log Kb= - log (Kw/Ka)

or, pKb = - log ( Kw/Ka)

Or, pKb = - log (1*10-14 / Ka )kindly put Ka value from question no. 6 . You get pKb value.